OpenStudy (anonymous):

A line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose all answers that are correct. A. −5x + y = −6 B. 5x + y = −6 C. (y-4)=-5(x+2) D. (y + 4) = −5(x − 2)

OpenStudy (igreen):

\(y - y_1 = m(x - x_1)\) Where y1 = y-value Where x1 = x-value Where m = slope

OpenStudy (compassionate):

A line has a slope of −5 and contains the point (−2, 4). We can use the formula y - y1 = m(x - x1) Where m = -5 -y1 = 4 and -x1 = -2

OpenStudy (anonymous):

So A and C?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Well, not exactly for A they're trying to trick you, because it's not in the correct form, you actually have to add 5 to both sides to get it into the correct form: \(-5x + y = 6\) Add 5x to both sides: \(y = 5x + 6\) So the slope of A, is 5, not -5, so it can't be an answer.

OpenStudy (igreen):

C is correct :)

OpenStudy (anonymous):

so B and C?

OpenStudy (igreen):

Yep@

OpenStudy (igreen):

*Yep!

OpenStudy (anonymous):

thank you!